OpenStudy (anonymous):

Area bounded between the curves: y=4cos(x) y=4cos(2x) Between 0

OpenStudy (anonymous):

Are you familiar with integrals?

OpenStudy (anonymous):

\[\int\limits_{0}^{\Pi}4\cos(2x)-4\cos(x)dx\]

OpenStudy (anonymous):

Do you need help solving this?

OpenStudy (anonymous):

For the most part! I know what to do; i had a problem with changing the limits of integration---- but turns out you don't even need to use u substitution. I'ma try it one more time--- if not. i'll bump it up!

OpenStudy (anonymous):

Okay. :)

OpenStudy (anonymous):

Hey, wanted to say thank you!